Use the following demand-and-cost information vis-à-vis three firms: Firm A, Firm B, and Firm C operating in three different market structures in the short run, to answer the questions that follow.


Firm A Firm B Firm C

Price where output is equal to zero 2000 1500 60

Profit maximizing price (R) 1000 ? 60

Profit maximizing output level (units) 30 50 2500

MC at profit maximizing output level (R) ? 275 60

ATC at profit-maximizing output level (R) 1300 ? 36

AVC at profit-maximizing output level (R) 900 300 ?

Minimum ATC (R) 1200 320 30

Minimum AVC (R) 800 280 3

Price at allocative efficient output level 160 580 ?

The allocative efficient level of output (units) 150 200 2500

Learner Index 0.8 ? 0

Total Fixed Cost (R) ? 2000 75000

Total Variable Cost at profit-maximizing output level 27000 15000 ?

Mark up ? 4 ?



4.1. Does Firm A, Firm B or Firm C operate in a perfectly competitive market?

Firm (A/B/C)______________

 

4.2. Calculate each of the following for Firm A:

MC = R______

Profit/loss = R_____

TFC = R_____

 


 

1
Expert's answer
2021-09-15T11:28:28-0400

a) From the given table we can understand that Firm C operates in a perfectly competitive market.

This is because in a perfectly competitive market, at the equilibrium, price = marginal cost of production. For firm C we see that profit-maximizing price is 60, which is same as the marginal cost.

Also, Lerner Index measures the degree of monopoly power of any firm. The value of this index is 0 for firm C thus implying that it has no monopoly power.

Answer: Firm C


b) For firm A :-

Lerner Index = 0.8

Now, Lerner Index "=\frac{ (P - MC)}{P}"

Hence

"\frac{ (P - MC)}{P}=0\\=\frac{ (1000- MC)}{1000}=0.8"

Solving,  MC = 200

"Profit/loss = price × output - total \space cost"

 

Now, total cost(TC) = Average total cost(ATC) × Output(Q)

Hence, "TC = 1300 \u00d7 30 = 39000"

 

Therefore, "Profit = 1000 \u00d7 30 - 39000 = - 9000"

So the firm is incurring a loss and loss = 9000


Total fixed cost(TFC)

= Average fixed cost(AFC) × Q

Now, "AFC = ATC - AVC"

         "= 1300 - 900\\

 = 400"

 

Hence, TFC = 400 × 30 = 12000